Minutes of the Tracker Meeting

September 30, 2004

 

Action Items:

- Arthur: write and send out a summary of the baseplate deformation issue.

- Arthur: contact QA people at Nyloc.

- Jeff: get the remaining studs dycronited asap.

- Jeff: procure a better socket wrench for stud installation.It needs to be longer with a tighter fit.

- Peter: get more extraction tools to Pisa for removing cones with the base plate in place.

- Dave: set up a meeting on the baseplate deformation issue.

- Sandro: update the 1584 procedure with references to the new procedures.

- Dave: set up a call-in for a G&A MRB at 9:00 am Wednesday.

- Luca: hand carry MCMs to Pisa Saturday (flight and non-flight).

†††† - Dave: get a high-level meeting going to push Parlex to perform.

††††† - Jeff: get 10 spare square nuts for test work at Nytemp.

††††† - Jeff: expedite an order for new cable clips for the cable holding fixture.

 

 

Agenda:

††††† - Status of tests on tray test articles.

††††† - Status of GSFC etching process.

††††† - Status of preparations for resuming flight production.

††††† - Status of alignment work.Discussion of issue with re-inserting studs.

††††† - Tower-interface nut/bolt design

††††† - MCM inspection and delivery for Tower A.

††††† - Status of other parts for Tower A:

††††††††††† - cables

††††††††††† - corner brackets

††††††††††† - shipping container

††††† - Tungsten reorder?

††††† - A.O.B.

 

Action Items from Last Week:

- Ben: weigh the tungsten foils before and after etch.

- Sandro: measure tile electrical isolation with a Keithley.

- Sandro: send a dimensioned sketch of his O-ring idea.

- Arthur: send photos of the cable clip fit problems.

- Martin: extract the clearance of the cable clip screw from the model.

- Sandro: send drawings next week of the ideas for resolving the connector clearance problem.

- Travelers to SLAC: bring shipping cases back.

- SLAC: ship cables with connector savers attached.

 

Ben sent the following information on the GSFC etching:

††††† We have completed testing of the T-peel and practice Bell peel tests made yesterday.The T-peel tests were done with 100-micron tiles with etches that removed 5% and 10% tungsten.All specimens were primed with BR-127 and bonded with 2216.Both etch types gave similar results: about 0.75 lb/in.

††††† We are now thinking that the tungsten is flexing during the T-peel and causing the BR-127 to separate.Note that BR-127 is more rigid than the 2216.The fact that it is very thin makes it act even more rigidly.This line of thinking would explain why the lap shear tests on pieces from the 700-micron showed the BR-127 with adhesion sufficient to result in 2216 cohesive failure (failure within the 2216), since the thick tungsten adherends did not flex during the lap shear tests.I am counting on seeing good results this afternoon when we test the thick and thin tungsten using a flatwise tension configuration, which again does not flex the tungsten.

††††† As an aside, the Plyform peel test (Bell peel) does not flex the tungsten either.Along these lines, I have Mike putting together Bell peel specimens in the Plyform configuration.Our practice tests showed the need to use a different adhesive, not EA 9394, to bond the aluminum adherends; and a need to bond the aluminum to the Kapton circuit sides separately before bonding the tungsten and Kapton with the 2216.

(Note added following the meeting: the flatwise tension test was highly successful.)

 

Sandro gave the status of the tray test articles in Rome.They have now been tested up to 70C.Small bubbles appear in all pieces, but in all cases they are smaller than the core cell size.The 3rd article (with no primer on the tungsten) has one complete delamination of Kapton from Tungsten (we don't yet know which side the glue remained on).Article 1 and 2 are okay at 70C.Article 1 had known trouble during the process of gluing on the top-side Kapton (the side without tungsten), and this face has problems not related to T/V test.The top sides of Articles 2 and 3 have only small bubbles that appeared at the higher temperatures.

 

Arthur reported that the alignment work has gone very well.However, there is an issue with the locking mechanism: the first test worked very well, with good running torque.But none of subsequent studs have behaved similarly, and they have had very high running torques. They were able to get 5 or 6 to work by running them in/out a few sides.The problem is on the Grid side as well as the nut side, even though with the EM test there was no problem on the Grid side.We need to talk to Nyloc about the tests they did and demand more testing.Also, Arthur proposed to do thorough tests at SLAC on 100% of the parts and match them with the grid simulator square nuts.Also, Arthur said that 1 square nut had some thread damage but was okay after running a stud in and out.36 studs are presently in Pisa; one or two have been lost.

Arthur asked Jeff to get the remaining studs dycronited.

He also said that we need custom tooling for a longer, better fitting socket for driving the studs.

Arthur said that one inner cone came out of the outer cone when the stud was removed.They dithered it back in place and kept the alignment intact, so that actually was a good sign.

Pisa needs more extraction tools for when the base plate is present.Peter is making more.

Arthur still needs to take photos of the clip issues.

 

There is an issue with the alignment work.Repeatability of measurements of the Grid simulator are not good to better 130 microns, due to changes with how it is held or clamped.Arthur also said that if you remove a stud and immediately replace it, it is fine.But if you take a stud out and put some weight on the Grid simulator, then the position is lost and the stud won't go back in.In general there is a feeling in Pisa that the grid simulator is not stiff enough.However, there really isnít space in any dimension to make it larger, and Sandro expressed a strong desire not to make it significantly heavier and hard to work with.We need a meeting of Ruland, Rich, Nordby, Brez etc. to discuss this.

 

Sandro is preparing summary of the test article work.The flight production baseline will be etched, primed tiles and curing at 35C.No grounding of tiles will be done.We will rely on the fact that the resistivity of the glue is such that the time constant for discharge is the order of a second, in which case there is no way the tiles could be charged up by cosmic-ray flux.

The Plyform tent will be installed by Monday.INFN will pick up etched tiles on Monday from the Torino company.All other materials are ready at INFN or Plyform.Sandro will update the 1584 procedure with references to the relevant procedures.16 bare trays were done by yesterday.Plyform is building a bottom tray today.20 bare trays will be ready by the weekend.

The MRR is scheduled for 8:00 am tomorrow.

 

We decided to have an MRB for the G&A production issues 9:00 am Wednesday.The topics are how to validate trays with no-encapsulation from MCM to SSD and the G&A solution to keep connectors below the surface.SLAC needs to modify the tray assembly drawings to remove encapsulation and to correct the tolerances for MCM installation.Sandro sent out a Word file describing the proposed solution.

SLAC needs redlines for the assembly drawing tolerances from Sandro.

 

Inspection of conformal coating of MCMs is starting today.The pitch adapter traces have been inspected.Marcus will execute electrical tests to weed out MCMs with poor margin with respect to clock duty cycle.Jeff will support the shipping.The goal is for Luca Latronico to hand carry the MCMs on Saturday.

 

SLAC will send 3 nonflight MCMs to support the cable testing.

Sandro wants 6 MCMs at least to work with at G&A in fixing the MCM bonding procedure.SLAC will send 6 functional preproduction units (Version 6 GTRC) for this purpose.There is no stock of nonfunctional MCMs (we generally donít accept them from Teledyne).

 

Arthur suggested making another cable bending tool in the Pisa shop.He is due back at SLAC Tuesday the 12th.

 

Corner brackets: new ones will be alodyned.They are now going into the plating shop.We should have them at SLAC next week.There is still a need to resolve the clip fit issues.Arthur said that the top bend will be done Monday on Tower-0, so will know more then.So far we only have info from the mockup.

 

Dave reported on the flex-circuit cable procurement.His blood is boiling, and he needs a meeting between Lowell and the CEO of Parlex; he is planning to threaten to cancel payment to Parlex of the expediting fee in the procurement contract (about $160,000).

 

Jeff will check the ETA of the outer shipping container.It is in fabrication.

Arthur wants Jeff to get 10 spare square nuts to send to Nytemp.

 

Dave wants to send another cable holding plate and clips to Pisa.Jeff will expedite new clips with the ESD cap on a credit card order.